OpenStudy (anonymous):

what is the slope of the lines through (-4, 2) and (5,8) ?

OpenStudy (anonymous):

Using the formula: rise over run. Slope = (y2-y1) / (x2-x1) slope = (8 - 2) / (5 - - 4) slope = 6 / 9 slope = 2/3

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (anonymous):

welcome